STATE-SPECIFIC DOUBLETS IN THE FTIR SPECTRUM OF GASEOUS TROPOLONE

\underline{R. L. REDINGTON}, Department of Chemistry and Biochemistry, Texas Tech University, Lubbock, TX 79409; R. L. SAMS, Environmental Molecular Sciences Laboratory, Pacific Northwest National Laboratory, Richland, WA 99352.